# The Gorini-Kossakowski-Sudarshan-Lindblad problem and the geometry of CP maps

### v3

@@ 11.3.3 Individual Lindblad parametrizations
Given $\Lambda\colon\mathsf{HP}(\mathcal{H})$, every $d$-cluster-point of $(\Delta_{\alpha}\Lambda)$, of which there is at least one, is a Lindblad parametrization of $\Lambda$.
@@
$(\Delta_{\alpha}\Lambda)$ is bounded by Lemma 11.1, hence has $d$-cluster-points (§9). Let $Z$ be one. By Lemma 11.3, LZLZ is a $d$-cluster point of the sequence $(L(\Delta_{\alpha}\Lambda))$. But Lemma 11.2 says that only $\Lambda$ can be such a cluster point. ∎
@@ 11.3.4 Full Lindblad parametrizer
In $\mathsf{Lind\,{\mathcal{H}}}$ equipped with the $d$-metric, $\left\{\Delta_{\alpha}\Lambda_{n}\,\colon\,\alpha\in{\mathbb{N}}\right\}$ is a precompact set. The product space $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$ can be turned into a metric space just as in §9, such that products of compact sets are compact. Thus, the sequence $\alpha\rightarrow\prod_{n}\Delta_{\alpha}\Lambda_{n}$ is precompact in $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$. $L^{\mathbb{N}}$ carries any cluster point of it into a cluster point of the sequence $\alpha\rightarrow\prod_{n}L(\Delta_{\alpha}\Lambda_{n})$ in $\mathsf{HP}(\mathcal{H})^{\mathbb{N}}$, by Lemma 11.3.
Thus, there is some subsequence, which we will just call $\Delta_{\beta}$ for simplicity, and a (not yet known to be linear) function $\Delta$ such that $\Delta_{\beta}\Lambda_{n}\xrightarrow{d}\Delta\Lambda$ and $(L\circ\Delta)\Lambda_{n}=\Lambda_{n}$ (Lemma 11.2) for each $n$. By $\|\Delta_{\beta}\|\leq M$ and Lemma 7.1(c), $\Delta_{\beta}\xrightarrow{SOT}\Delta$ and $\Delta$ is linear with $\|\Delta\|\leq M$. Finally, by boundedness of $L$, $L\circ\Delta=\operatorname{\mathbbm{1}}_{\mathsf{HP}(\mathcal{H})}$.
That $\Delta$ maps $\mathsf{cp}^{+}(\mathcal{H})$ into $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ follows just as in Lemma 11.4. ∎
@@ 12 Conclusion
We have shown that provides an excellent foundation on which to build the basic finite-dimensional theory of CP maps and of generators of Markovian CP evolutions in a unified way. For a small upfront investment, everything flows very smoothly. All the basic results can be extended to the separable setting by use of finite-dimensional approximations in the form of filtrations. It is significant that this can be done, because there is an enormous gulf between the conventional approaches to the finite-dimensional and separable cases.
@@ Appendix A Glossary/Index
$\square\,$ Def. 2.1
$\bm{\pi}$} Def. 2.2
$\bm{\theta}$} Def. 3.1

### v4

@@ 11.3.4 Individual Lindblad parametrizations
Given $\Lambda\colon\mathsf{HP}(\mathcal{H})$, every $d$-cluster-point of $(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)$, of which there is at least one, is a Lindblad parametrization of $\Lambda$.
@@
$(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)$ is bounded by Lemma 11.1, hence has $d$-cluster-points (see §9). Let $Z$ be one such. By Lemma 11.4, $LZ$ is a $d$-cluster point of the sequence $(L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda))$. But Lemma 11.3 says that only $\Lambda$ can be such a cluster point.
Finally, if $\Lambda$ is $\mathsf{cp}^{+}$, then the first component of $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP, according to Lemma 11.2. Since $\mathsf{CP}(\mathcal{H})$ is $d$-closed (Prop. 10.5), every $d$-cluster point of $(\Delta^{\scriptstyle{min}}_{\alpha})$ also has this property. ∎
@@ 11.3.5 Full Lindblad parametrizer
In $\mathsf{Lind\,{\mathcal{H}}}$ equipped with the $d$-metric, $\left\{\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}\,\colon\,\alpha\in{\mathbb{N}}\right\}$ is a precompact set. The product space $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$ can be turned into a metric space just as in §9, such that products of compact sets are compact. Thus, the sequence $\alpha\mapsto\prod_{n}\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}$ is precompact in $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$. $L^{\mathbb{N}}$ carries any cluster point of it into a cluster point of the sequence $\alpha\mapsto\prod_{n}L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n})$ in $\mathsf{HP}(\mathcal{H})^{\mathbb{N}}$, by Lemma 11.4.
Thus, there is some subsequence, which we will just call $\Delta^{\scriptstyle{min}}_{\beta}$ for simplicity, and a (not yet known to be linear) function $\Delta$ such that $\Delta^{\scriptstyle{min}}_{\beta}\Lambda_{n}\xrightarrow{d}\Delta\Lambda$ and $(L\circ\Delta)\Lambda_{n}=\Lambda_{n}$ (Lemma 11.3) for each $n$. By $\|\Delta^{\scriptstyle{min}}_{\beta}\|\leq M$ and Lemma 7.1(c), $\Delta^{\scriptstyle{min}}_{\beta}\xrightarrow{SOT}\Delta$ and $\Delta$ is linear with $\|\Delta\|\leq M$. Finally, by boundedness of $L$, $L\circ\Delta=\operatorname{\mathbbm{1}}_{\mathsf{HP}(\mathcal{H})}$.
That $\Delta$ maps $\mathsf{cp}^{+}(\mathcal{H})$ into $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ follows just as in Lemma 11.5. ∎
11.4 The tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ and tangent space $\mathsf{cp}(\mathcal{H})$ are $d$-closed
In §10.2, we showed that $\mathsf{CP}(\mathcal{H})$ is $d$-closed. Now we will show the same holds for $\mathsf{cp}^{+}(\mathcal{H})$, and therefore for $\mathsf{cp}(\mathcal{H})$ as well, since $\mathsf{cp}(\mathcal{H})=\mathsf{cp}^{+}(\mathcal{H})\cap(-\mathsf{cp}^{+}(\mathcal{H}))$.
To that end, suppose $(\Lambda_{n})$ is a sequence in $\mathsf{cp}^{+}(\mathcal{H})$, $d$-convergent to, say, $\Lambda$. In addition, assume that this sequence is bounded in norm. Later, we will see how to remove this restriction. By Prop. 11.6, or even Lemma 11.5, there is a bounded sequence $Z_{n}$ of Lindblad parametrizations in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. This latter sequence has $d$-cluster points; let $Z$ be one such. Necessarily, it too is in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. By Lemma 11.4, $LZ$ is a $d$-cluster point of $(\Lambda_{n})$. Since $\Lambda_{n}\xrightarrow{d}\Lambda$, it follows that $LZ=\Lambda$.
Returning to the deferred business, drop the assumption that $(\Lambda_{n})$ is bounded. We show how to construct a bounded sequence also converging to $\Lambda$. By definition of $d$-convergence, there is $N_{1}$ such that $n>N_{1}$ implies that $\|\mathscr{P}_{1}(\Lambda-\Lambda_{n})\|<1$, $N_{2}>N_{1}$ such that $n>N_{2}$ implies that $\|\mathscr{P}_{2}(\Lambda-\Lambda_{n})\|<1$, etc. Now, define $\Gamma_{n}=\mathscr{P}_{n}\Lambda_{N_{n}}$. The sequence $(\Gamma_{n})$ is bounded, and still $d$-converges to $\Lambda$.
@@ 12 Conclusion
We have shown that the Jamiołkowski transform $\mathfrak{J}$ provides an excellent foundation on which to build the basic finite-dimensional theory of CP maps and of generators of Markovian CP evolutions in a unified way, and that the basic results can be extended to the separable setting by use of finite-dimensional approximations in the form of filtrations and their associated metrics. Among these basic results are: extreme CP maps are $\bm{\theta}$}-maps, every CP map has Kraus decompositions, a time-dependent element of $\mathsf{cp}^{+}(\mathcal{H})$ which is norm continuous generates a Markovian CP evolution, and Lindblad parametrizers exist. In addition, the classes $\mathsf{CP}(\mathcal{H},\mathcal{K})$, $\mathsf{cp}^{+}(\mathcal{H})$ and $\mathsf{cp}(\mathcal{H})$ are $d$-closed. It is significant that this extension can be done for Kraus decompositions, in particular, because there is an enormous gulf between the conventional approaches to the finite-dimensional and separable cases. Perhaps not so much as regards the GKSL problem, since Lindblad [2] does, in fact, make use of a sequence of finite-dimensional approximations in his key Proposition 5. However, there seems to be considerable difference in the way that primitive idea is used there compared to this paper.
@@ Appendix A Glossary/Index
$\square\,$ $A\,\square\,B=\rho\mapsto A\rho B$, Def. 2.1
$\bm{\pi}$} ${\bm{\pi}}(h)=h\overline{h}$, Def. 2.2
$\bm{\theta}$} ${\bm{\theta}}(A)=A\,\square\,A^{\dagger}$, Def. 3.1